OpenStudy (anonymous):

You have two block of clay in cube form and the edges are 10 cm. How many spheres with a radius of 5 cm can you make with that amount of clay? help Me ..

OpenStudy (zehanz):

Divide the volume of the cube by the volume of the spheres: \(V_{cube}=10^3\) \(V_{sphere}=\frac{4}{3}\pi r^3=\frac{4}{3}\pi 5^3\) So calculate this: \(\dfrac{V_{cube}}{V_{sphere}}=\dfrac{1000}{\frac{4}{3}\cdot 125\pi}\) and then round down to get the number of spheres.